OpenStudy (anonymous):

need help on lim -->1 of sin(pi*x)/(x-1)

OpenStudy (experimentx):

use substitution x-1 = u, x= u+1 limit u -> 0 and see what happens

OpenStudy (anonymous):

hum.. let me see.

OpenStudy (jhannybean):

can't you use L'Hospital's rule,@experimentX since you have the indeterminant form 0/0?

OpenStudy (anonymous):

L'Hoptal's is not alowed.

OpenStudy (experimentx):

yes you can ... but not recommended if you can do algebraically. best substitution is to use x-1 = u/pi

OpenStudy (jhannybean):

:( ok.

OpenStudy (anonymous):

I get stucked on lim u-->0 of sin(u+1)/u

OpenStudy (experimentx):

no that is not correct form

OpenStudy (anonymous):

sin x/ x = 1 how to apply this?

OpenStudy (jhannybean):

replace the x in sin ( pi*x) with x = u+1

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I got lim u-->0 -sin(u*pi)/u

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

(-sin(u)/u)*sin(pi)?

OpenStudy (anonymous):

-1*sin(pi)?

OpenStudy (anonymous):

=0

OpenStudy (experimentx):

use another substitution \( \pi u = v \)

OpenStudy (anonymous):

ok, let me try...

OpenStudy (anonymous):

lim v->0 - sin(v)/v making v=pi*u = -1?

OpenStudy (experimentx):

change the lower u to v/pi

OpenStudy (experimentx):

you forgot to add pi to the top

OpenStudy (anonymous):

sorry, I'm lost, where i did the mistake?

OpenStudy (anonymous):

I forgot to add pi? where?

OpenStudy (experimentx):

\[ \frac{-\sin (\pi u)}{u} = ?\]

OpenStudy (anonymous):

.sin(pi*u+pi) = sin(pi*u)cos(pi) + sin(pi)cos(pi*u) sin (pi) = 0 and cos(pi) = -1 then -sin(pi*u)... where is the missing pi?

OpenStudy (experimentx):

put u = v/pi you get -sin(pi*v/pi)/(v/pi) = - pi sin(v)/v = -pi

OpenStudy (anonymous):

Nice. Thank you. But I need more exercicies like this. Can you indicate any link?

OpenStudy (experimentx):

don't know ... if you have used this x-1 = u/pi, you would have gotten easy answer. don't know ... try constructing few stuff yourself.

OpenStudy (anonymous):

ok. Thank you again!

OpenStudy (experimentx):

yw